OpenStudy (anonymous):

Create your own two-step equation with All Real Numbers as the solution. Show the steps to solving the equation to prove the solution All Real Numbers.

OpenStudy (ajanijones):



OpenStudy (phi):

Still working on this? an equation will all real numbers as the solution means an equation that is true no matter what number you pick. Example x= x where x is an "unknown" number. Notice that if x is 1, this equation becomes 1=1 which is true... meaning x=1 is a solution. Of course, it is true no matter what x you pick

OpenStudy (anonymous):

hahaha @phi yea i am lol

OpenStudy (phi):

You now have to make it more complicated start with x= x now multiply and add things to both sides to make it look more complicated.

OpenStudy (anonymous):

what "things"? numbers??

OpenStudy (phi):

like this: 3x = 3x (multiplied both sides by 3) 3x + 4 = 3x + 4 (added 4 to both sides) we could add letters: 3x + y + 4 = 3x + y + 4 (but we don't want it too complicated)

OpenStudy (anonymous):

ok i sort of get that.. how would it be all written out? in answer form?

OpenStudy (phi):

write down the equation. solve it.

OpenStudy (anonymous):

oh the equation u gave me?

OpenStudy (phi):

I would feel better if you picked your own equation.

OpenStudy (anonymous):

gah! ok lol :P *inner turmoil* >.<

OpenStudy (anonymous):

this is so confusing. :/

OpenStudy (phi):

start with x= x now multiply and add things to both sides to make it look more complicated.

OpenStudy (phi):

start with x=x pick a number, and multiply both sides by that number.

OpenStudy (anonymous):

ok. im picking a number

OpenStudy (anonymous):

5x=5x?

OpenStudy (anonymous):

multipy both sides by 5?

OpenStudy (phi):

now add the same number to both sides (they want a 2-step equation)

OpenStudy (anonymous):

5x + 2= 5x+2 ?

OpenStudy (anonymous):

where does all real numbers come in? i dont understand what thats supposed to mean

OpenStudy (phi):

yes, that is an equation. now "solve it" which means find all the x values that make that equation true.

OpenStudy (anonymous):

oooooooh ok i didnt get that all this time

OpenStudy (anonymous):

but now how do i solve it haha??

OpenStudy (phi):

to solve equations, you simplify them. For example add -2 to both sides, like this 5x + 2 -2 = 5x + 2 - 2 now simplify 2-2 is 0 5x+0 = 5x +0 and 5x+0 is 5x (0 added to anything does not change it) now you have 5x= 5x

OpenStudy (phi):

next step: add -5x to both sids

OpenStudy (anonymous):

oh like when ur supposed to do the opposite ... to get the answer

OpenStudy (phi):

yes. but in this case you will end up with 0 = 0 that means the original equation is true for any x value. In other words, it is true for all real numbers.

OpenStudy (anonymous):

ok thanks! i figured it out :))

OpenStudy (phi):

Hopefully your lesson explains this somewheres.

OpenStudy (anonymous):

haha but my brain doesnt process it 99.99 % of the time :P

OpenStudy (phi):

Most of the ideas are not hard... just different so you never thought of them before. If you get stuck, stop and say to yourself "this does make sense... what idea did I miss?" and ask questions. and don't all asleep.

OpenStudy (anonymous):

haha ok thanks